A thermally insulated water heater contains 83 liters of water. How much electric energy in kWh is required to increase the water temperature 21ºC.
Answer #1

Electrical Energy required required to raise the water temperature to 21 degree cel is

P =(4.2 * T * L) / (3600)

T= temperature difference

L=number of liters of water to be heated

P= (4.2 * 21 * 83 ) / 3600

P = (7320.6) / 3600

P = 2.0335 kWh